| Type | Comment | count | priority | fname | location | serial | time | value | min | max |
|---|---|---|---|---|---|---|---|---|---|---|
| hg: | Lamp not detected f1= 11 f4= 15 | 2 | 1 | B30022.194 | Toron | #194 | 11:03:23 | 1 | -1 | |
| hg: | Lamp not on... hg test terminated F1= 11 F4= 15 | 2 | 1 | B30022.194 | Toron | #194 | 11:03:23 | 1 | -1 | |
| hp: | Spectrometer alignment trouble. | 2 | 1 | B30022.194 | Toron | #194 | 12:33:48 | 1 | -1 | |
| hg: | No intensity detected. Check PMT. | 1 | 1 | B30022.194 | Toron | #194 | 12:36:30 | 1 | -1 |
| Serial | YYYYJJJ | HG large | max correction |
|---|---|---|---|
| 194 | 2022300 | 273 | 7.1 |
| Serial | YYYYJJJ | HP large | max correction |
|---|---|---|---|
| 194 | 2022300 | 5 | 72.5 |
| Type | Comment | count | priority | fname | location | serial | time | value | min | max |
|---|---|---|---|---|---|---|---|---|---|---|
| fr: | Micrometer #1 reference found at step 9905 (+CAL)= 10908 Now set to o3 offset = 7985 | 1 | 99 | B30022.194 | Toron | #194 | 12:29:19 | 10908 | 1000 | 9000 |
| dsp: | Doing dispersion test with Hg (UV port) | 3 | 99 | B30022.194 | Toron | #194 | 15:08:25 | 1 | -1 | |
| checkhg: | Last HG had to adjust the micrometer by at least 1 | 3 | 99 | B30022.194 | Toron | #194 | 15:50:26 | 1 | -1 |
| _jday | _loc | _serial | _skc | dt | hg | hp | rs | sl |
|---|---|---|---|---|---|---|---|---|
| 300 | Toronto | #194 | menu,Must do HG now,Starting schedule inside | 2 | 41 | 45 | 2 | 2 |
| Serial | RH, % | AH, g/m3 | 7-day RH history | Des-nt replcd days ago |
|---|---|---|---|---|
| 194 | 25.8 | 5.3 |